Molteni Motors Inc. recently reported $2.25 million of net income. Its EBIT was $7 million, and its tax rate was 40%. What was its interest expense? Round your answer to the nearest dollar. Enter your answer in dollars. For example, an answer of $1.2 million should be entered as 1,200,000.
Answer #1

Net income = 2.25 million
EBIT = 7 million
Tax rate = 40%

(EBIT - Interest )* (1-Tax Rate) = Net Income
(7 - Interest ) *(1-40%) = 2.25
7-Interest = 2.25/0.6
Interest =7 -2.25/0.6 =3.25 million or $3,250,000